OpenStudy (anonymous):

Rewrite the function y=5+15x/5x in simplest terms.

OpenStudy (anonymous):

\[\frac{5+15x}{5x}\] like that?

OpenStudy (anonymous):

yeah!

OpenStudy (anonymous):

cancel a 5 top and bottom

OpenStudy (australopithecus):

well notice that (x+y)/x = x/x + y/x

OpenStudy (anonymous):

\[\frac{5+15x}{5x}=\frac{5(1+3x}{5x}=\frac{1+3x}{x}\]

OpenStudy (australopithecus):

so 5/5x + 15x/5x = 1/x + 15/5 = 1/x + 3

OpenStudy (australopithecus):

or (1 + 3x)/x

OpenStudy (anonymous):

also what austra said if you wan to break it apart